Power Manufacturing has equipment that it purchased 7 years ago for $2,750,000. The equipment was used for a project that was intended to last for 9 years. However, due to low demand, the project is being shut down. The equipment was depreciated using the straight-line method and can be sold for $440,000 today. The company's tax rate is 34 percent. What is the aftertax salvage value of the equipment?

  • $440,000

  • $381,822

  • $440,000

  • $498,178

  • $589,600
Answer #1

Annual depreciation=(Cost-Salvage value)/Useful Life

=(2,750,000/9)=$305,555.56/year(Approx)

Hence book value as on date of sale=2,750,000-(305,555.56*7)

=$611,111.11

Hence loss on sale =$611,111.11-$440,000=$171,111.11

After-tax salvage value=Sale proceeds+(Tax rate*loss on sale)

=440,000+(171,111.11*0.34)

=$498178(Approx).